OpenStudy (anonymous):

X^2-x=2 I know x=2 but i need the work

OpenStudy (anonymous):

\[x^{2} - x - 2 = 0\]

ganeshie8 (ganeshie8):

how about x=-1, it works too right ?

ganeshie8 (ganeshie8):

in general, if u have degree of 2, you will have 2 solutions that will work

OpenStudy (anonymous):

you can factor and set each factor = 0 and solve for x.

OpenStudy (anonymous):

use the quadratic formula on your second order polynomial in the form of ax^2 + bx + c = 0 \[x_{1and2} = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a}\]

OpenStudy (anonymous):

1st x from using the minus sign in plus-minus and 2nd x using the plus sign

OpenStudy (anonymous):

\[x^2 - x - 2 = \left(x-2\right)\left(x+1\right)=0 \Rightarrow x-2 = 0 \text{ or } x+1 = 0\]